Twenty-one more than three times a number is twelve less than the sum of the number and 5. Find the number.
Answer by nerdybill(7384)

Twenty-one more than three times a number is twelve less than the sum of the number and 5. Find the number.
.
Let x = "the number"
then
3x + 21 = (x+5) - 12
3x + 21 = x - 7
2x + 21 = -7
2x = -28
x = -14
.
The number is -14